The haloween association reported that last year Americans spent $0.68 billion more on candy than costumes for Halloween. If the total spent by Americans for both items was 3.18 billion, how much did Americans spend on each item?
Answer by checkley79(3341)

X+(X+.68)=3.18
2X=3.18-.68
2X=2.5
X=2.5/2
X=1.25 BILLION FOR THE HALLOWEEN COSTUMES.
3.18-1.25=1.93 BILLION FOR CANDY.